OpenStudy (anonymous):

Solve the following compound inequality. Describe the solution set using interval notation.

OpenStudy (anonymous):

\[\frac{1}{3}< \frac{7}{6}(l-3)< \frac{2}{3}\]

OpenStudy (anonymous):